# Little Red Dots are Tidal Disruption Events in Runaway-Collapsing Clusters

### v1

@@ 1 Introduction
The objects known as Little Red Dots (LRDs) discovered by the James Webb Space Telescope (JWST) have been mystifying our community (Greene et al., 2024; Kocevski et al., 2024; Kokorev et al., 2024; Matthee et al., 2024). They exhibit features common to both star-forming galaxies and active galactic nuclei (AGNs), but their emission cannot be modeled robustly by either (or a combination). Their number density at high redshift is also higher than expected compared to similar galaxy populations. If they are dominated by star formation, they are more compact than anything ever seen before (Guia et al., 2024a), and exceed the predicted stellar mass function allowed by CDM (Boylan-Kolchin, 2023; Baggen et al., 2024a; Akins et al., 2024). If they are dominated by AGN emission, their existence implies an almost 100% AGN fraction in high-redshift galaxies (Greene et al., 2024). Both of these possibilities are difficult to reconcile with any cosmic evolution model.
@@ 2 Runaway Collapse
One prominent theory for the formation of black hole seeds is the runaway collapse of dense star clusters. It is thought to create seeds of and may or may not be redshift dependent, depending on whether metallicity is an important factor. The phenomenon of runaway collapse involves a dynamical instability resulting in the collapse of the cluster core. This process may result in a supermassive star, which then evolves into an intermediate mass black hole (Devecchi & Volonteri, 2009), or may involve a pre-existing black hole which rapidly accretes surrounding objects in the cluster (Guia et al., 2024b; Baggen et al., 2024b; Devecchi & Volonteri, 2009; Davies et al., 2011). Regardless of the precise order of operations, a dense cluster surrounding a massive black hole will be an efficient environment for tidal disruption events. In fact, Alexander & Bar-Or (2017) suggest that intermediate mass black holes should not exist at all because growth from TDEs is so efficient that the minimum MBH mass is .
There have been very few calculations of the number density of such seeds with cosmic time, likely because the conditions for runaway collapse happen on a much smaller scale than that of e.g. cosmological simulations. Estimates have been made by Devecchi et al. (2012) (analytic) and Habouzit et al. (2017) (numerical) which I show in Figure 1. The grey shaded area represents an adaptation of the prediction from Devecchi et al. (2012) for a range of seed masses (). The solid lines are from a cosmological hydrodynamic simulation with a using a subgrid model for cluster-based IMBH formation (Habouzit et al., 2017). The blue line shows a model with delayed cooling after supernova explosions, while the orange represents regular kinetic or thermal supernova feedback.
Also included in this figure are observational results of LRDs from JWST: Greene et al. (2024) (green dots), Kocevski et al. (2024) (red stars) and Kokorev et al. (2024) (black diamonds). The observed points are about 4 orders of magnitude below the predicted density of black hole seeds formed through cluster collapse. A cannonical TDE has a peak bolometric luminosity of erg s-1 and remains observable for about 1 year (Gezari, 2021). Therefore, a TDE rate of per year could explain the fraction of observed LRDs vs the occupation of black holes in galaxies.
@@ 3 TDE Rates in LRDs
Here is the total mass (), and is the tidal radius given by and is the solar radius multiplied by .
@@
where is a numerical prefactor set to 0.8 and is the fraction of bound stars, which is assumed to be 0.2. Using these two rate estimates, I compute expected TDE rates in LRD environments. The stellar densities in LRDs are estimated to be in the realm of pc-3, and I use as a fiducial value (Guia et al., 2024a). Using a canonical velocity dispersion for cluster collapse estimated at 40 km s-1 (Miller & Davies, 2012) and a mean stellar mass of 1 M⊙, I plot the TDE rate in Figure 2 as a function of black hole mass (top) and stellar velocity dispersion (bottom). In the top plot the velocity dispersion is fixed at 40 km s-1, while the bottom plot has a fixed M M⊙. The rate estimate by Stone et al. (2017) is plotted in blue and Rizzuto et al. (2023) in red. A rate of per year is very reasonable within the range of and/or velocity dispersions of 100 - 150 km s-1. There are no direct constraints on stellar velocity dispersions in LRDs, but it is plausible that with such high densities the values would be large.
@@ 4 Conclusions and Caveats
These issues can be avoided if one assumes a non-standard initial mass function (IMF). A top-heavy IMF would result in higher luminosity stars, and an overall decrease in stellar mass estimates. I propose that the source of the light is stellar, originating in the dense nuclear cluster surrounding the IMBH, and with a top-heavy IMF. The whole issue of too bright / too many stars is solved if there are actually fewer stars of higher masses. A top-heavy IMF from a metal-poor population could produce the red stellar light, explain a lack of dust (if the environment is very metal-poor), and solve the number density problem. More massive stars would also result in more luminous TDEs. SED models of LRDs with varying IMFs are needed to confirm this suggestion.
LRDs are not all identical, and there is no need for one overarching theory to explain their existence. For example, some objects show strong AGN signatures (Labbe et al., 2024; Wang et al., 2024, see) which likely dominate the SEDs. Therefore, TDEs occurring in runaway collapsing clusters may explain a subset of LRDs, and canonical AGN and other phenomena the others. One way to verify the TDE hypothesis is to search for strong variability in the rest-frame UV (caused by TDE emission) but less in the rest-frame optical/NIR (dominated by stellar light). TDEs have well-known decays in their light curves which can be confirmed fairly straightforwardly. The caveat is that due to time dilation timescales are lengthened by a factor of , so decays will occur approximately 6-8 times more slowly than seen in the local universe. Initial searches for variability of LRDs have found none as of yet (Kokubo & Harikane, 2024; Tee et al., 2024), highlighting that these objects may have a diverse taxonomy. Follow-up with JWST at appropriate timescales is crucial for verifying the TDE hypothesis.

### v2

@@ 1 Introduction
The objects known as Little Red Dots (LRDs) discovered by the James Webb Space Telescope (JWST) have been mystifying our community (Greene et al., 2024; Kocevski et al., 2024; Kokorev et al., 2024; Matthee et al., 2024). They exhibit features common to both star-forming galaxies and active galactic nuclei (AGNs), but their emission cannot be modeled robustly by either (or a combination). Their number density at high redshift is also higher than expected compared to similar galaxy populations. If they are dominated by star formation, they are more compact than anything ever seen before (Guia et al., 2024), and exceed the predicted stellar mass function allowed by CDM (Boylan-Kolchin, 2023; Baggen et al., 2024; Akins et al., 2024). If they are dominated by AGN emission, their existence implies an almost 100% AGN fraction in high-redshift galaxies (Greene et al., 2024). Both of these possibilities are difficult to reconcile with any cosmic evolution model.
@@ 2 Runaway Collapse
One prominent theory for the formation of black hole seeds is the runaway collapse of dense star clusters. It is thought to create seeds of and may or may not be redshift dependent, depending on whether metallicity is an important factor. The phenomenon of runaway collapse involves a dynamical instability resulting in the collapse of the cluster core. The ensuing runaway collisions in these nuclear star clusters may result in the creation of a supermassive star (Portegies Zwart et al., 1999; Ebisuzaki et al., 2001; Portegies Zwart et al., 2004; Devecchi & Volonteri, 2009) which then evolves into an intermediate mass black hole. This phenomenon has been demonstrated in Monte Carlo (Gürkan et al., 2004; Freitag et al., 2006) and N-body (Rantala & Naab, 2025) simulations, and runaway collapse may be enhanced in cluster with low metallicities and higher masses (Di Carlo et al., 2021) and/or high binary fractions of massive stars (González et al., 2021). Other works suggest IMBH formation may involve a pre-existing black hole which rapidly accretes surrounding objects in the cluster (Begelman & Rees, 1978; Davies et al., 2011). Regardless of the precise order of operations, a dense cluster surrounding a massive black hole will be an efficient environment for tidal disruption events. In fact, Alexander & Bar-Or (2017) suggest that intermediate mass black holes should not exist at all because growth from TDEs is so efficient that the minimum MBH mass is .
There have been very few calculations of the number density of such seeds with cosmic time, likely because the conditions for runaway collapse happen on a much smaller scale than that of e.g. cosmological simulations. Estimates have been made by Devecchi et al. (2012) (analytic) and Habouzit et al. (2017) (numerical) which I show in Figure 1. The grey shaded area represents an adaptation of the prediction from Devecchi et al. (2012) for a range of seed masses (). The solid lines are from a cosmological hydrodynamic simulation using a subgrid model with star-formation-based IMBH formation (Habouzit et al., 2017). The blue line shows a model with delayed cooling after supernova explosions, while the orange represents regular kinetic or thermal supernova feedback. Their model includes a combination of a cluster-based model, which is based on the formation of a supermassive star as described in the previous paragraph, and seeds from Population III stars. This model specifically aims to create seeds based on the formation of the first stars and nuclear star clusters. Both types of seeds are formed by finding dense pristine clumps of gas with an average mass of M⊙. Their results do not differentiate the two threads of seed formation, so the data in Figure 1 may be taken as an upper limit. Also included in this figure are observational results of LRDs from JWST: Greene et al. (2024) (green dots), Kocevski et al. (2024) (red stars) and Kokorev et al. (2024) (black diamonds). The observed points are about 4 orders of magnitude below the predicted density of black hole seeds formed through cluster collapse. A cannonical TDE has a peak bolometric luminosity of erg s-1 and remains observable for about 1 year (Gezari, 2021). Therefore, a TDE rate of per year could explain the fraction of observed LRDs vs the occupation of black holes in galaxies.
@@ 3 TDE Rates in LRDs
Here is the total mass (), and is the tidal radius given by and is the solar radius multiplied by . Tidal capture is likely dominant at lower black hole masses (e.g. ) and would contribute minimal luminosity. A possible limitation of this model is that initial black hole growth may be delayed because tidally captured stars will inflate before being consumed, causing a slow build before runaway growth occurs. Therefore, it is uncertain whether this method will be fast enough to jumpstart the growth of LRD black holes by the time they are observed. As the dominant process shifts to TDEs at higher masses, Stone et al. (2017) argue that in the full loss cone approximation, the growth rate would remain similar and still scale as . However, it is also unclear at which point, if any, the growth of the black hole transitions from being dominated by tidal captures to tidal disruption events. The evolution of a collapsing cluster is dynamic, and may include further gas inflow from the larger environment (and subsequent star formation), so it is likely that a full loss cone is an appropriate assumption.
@@
where is a numerical prefactor set to 0.8 and is the fraction of bound stars, which is assumed to be 0.2. This analytic rate matches their simulations well, and while the simulations were only run for black hole masses growth up to for computational reasons, they suggest the mass growth will scale linearly with the TDE rate, which will remain as calculated above as long as the physical conditions (e.g. density, velocity dispersion) of the cluster are unchanging. These simulations assume old stellar ages and therefore do not include effects from supernovae or other stellar feedback in the cluster. They also do not include the effect of multiple black holes, which could increase (by increasing the cross-section) or decrease (by scattering objects out of the cluster) the TDE rate. Overall, both rate estimates come from different rationales and are in rough agreement with each other, so they can be used as a guideline.
Using these two rate estimates, I compute expected TDE rates in LRD environments. The stellar densities in LRDs are estimated to be in the realm of pc-3, and I use as a fiducial value (Guia et al., 2024). Using a canonical velocity dispersion for cluster collapse estimated at 40 km s-1 (Miller & Davies, 2012) and a mean stellar mass of 1 M⊙, I plot the TDE rate in Figure 2 as a function of black hole mass (top) and stellar velocity dispersion (bottom) (thick lines). In the top plot the velocity dispersion is fixed at 40 km s-1, while the bottom plot has a fixed M M⊙. The rate estimate by Stone et al. (2017) is plotted in blue and Rizzuto et al. (2023) in red. A rate of per year is very reasonable within the range of and/or velocity dispersions of 100 - 150 km s-1. There are no direct constraints on stellar velocity dispersions in LRDs, but it is plausible that with such high densities the values would be large.
Shown in the dashed, dotted, and dot-dashed lines in Figure 2 are rates calculated bracketing a reasonable LRD parameter space. The dotted lines display a fiducial stellar density of pc-3 rather than , and show that for the lower bound of estimated densities the TDE rate may still reach year-1 but only for larger black holes and lower velocity dispersions. The dashed and dot-dashed lines show extrema in velocity dispersion (150 km s-1, upper panel) and black hole mass ( and M⊙, respectively (lower panel).
As mentioned above, the tidal capture prediction of Stone et al. (2017) is only valid for the full loss-cone assumption. At some critical mass the loss cone transitions from full to empty, at which point the rate of TDEs will scale as M. We include this calculation in the form of light grey downward-sloping lines in Figure 2 for a range of values. (We include them on the fiducial calculation only so as not to further distract from the results of the figure.) In many cases the TDE rate increases with increasing black hole mass and decreasing velocity dispersion, and a rate of year-1 is unlikely to be reached in a case of small black hole mass and/or high velocity dispersion.
@@ 4 Conclusions and Caveats
These issues can be avoided if one assumes a non-standard initial mass function (IMF). A top-heavy IMF would result in higher luminosity stars, and an overall decrease in stellar mass estimates. The IMF has been shown to vary with density and metallicity (Marks et al., 2012), which are both extreme in the case of LRDs. Jeřábková et al. (2017) has shown that star clusters with top-heavy IMFs can reach up to quasar luminosities (i.e. erg s-1 for a stellar mass of ), with a mass to light ratio of , for the first 10 Myr of their lifetimes. Therefore, I propose that the source of the light is stellar, originating in the dense nuclear cluster surrounding the IMBH, and with a top-heavy IMF. The whole issue of too bright / too many stars is solved if there are actually fewer stars of higher masses. A top-heavy IMF from a metal-poor population could sufficiently produce the red stellar light, explain a lack of dust (if the environment is very metal-poor), and solve the number density problem. More massive stars would also result in more luminous TDEs. SED models of LRDs with varying IMFs are needed to confirm this suggestion.
LRDs are not all identical, and there is no need for one overarching theory to explain their existence. For example, some objects show strong AGN signatures (see Labbe et al., 2024; Wang et al., 2024) which likely dominate the SEDs. Therefore, TDEs occurring in runaway collapsing clusters may explain a subset of LRDs, and canonical AGN and other phenomena the others. One way to verify the TDE hypothesis is to search for strong variability in the rest-frame UV (caused by TDE emission) but less in the rest-frame optical/NIR (dominated by stellar light). TDEs have well-known decays in their light curves which can be confirmed fairly straightforwardly. TDEs have also been shown to exhibit decaying luminosity in coronal lines on timescales of years (Wang et al., 2012), which could be observed with spectroscopic follow-up. The caveat is that due to time dilation timescales are lengthened by a factor of , so decays will occur approximately 6-8 times more slowly than seen in the local universe. Initial searches for variability of LRDs have found none as of yet (Kokubo & Harikane, 2024; Tee et al., 2024), highlighting that these objects may have a diverse taxonomy. Follow-up with JWST at appropriate timescales is crucial for verifying the TDE hypothesis.